# Metastability and Ostwald Step Rule in the Crystallisation of Diamond and Graphite from Molten Carbon

### v1

@@
Metastability and Ostwald Step Rule in the Crystallisation of Diamond and Graphite from Molten Carbon
Davide Donadio Department of Chemistry, University of California Davis ddonadio@ucdavis.edu.    Margaret L. Berrens Department of Chemistry, University of California Davis Quantum Simulations Group, Physics Division, Lawrence Livermore National Laboratory, Livermore, California 94550, United States.    Wanyu Zhao    Shunda Chen    Tianshu Li Department of Civil and Environmental Engineering, George Washington University
(March 3, 2025)
@@
The crystallisation of carbon from the melt under extreme conditions is highly relevant to earth and planetary science,Timmerman et al. (2023); Ross (1981) materials manufacturing, and nuclear fusion research.Zylstra et al. (2022) The thermodynamic conditions near the graphite-diamond-liquid (GDL) triple point are especially of interest for geological and technological applications, but high-pressure flash heating experiments aiming to resolve this region of the phase diagram of carbon exhibit large discrepancies.Bundy et al. (1996); Togaya (1997); Zhao et al. (2013); Yang et al. (2023) Experimental challenges are often related to the persistence of metastable crystalline or glassy phases, superheated crystals, or supercooled liquids. A deeper understanding of the crystallisation kinetics of diamond and graphite is crucial for effectively interpreting the outcomes of these experiments. Here, we reveal the microscopic mechanisms of diamond and graphite nucleation from liquid carbon through molecular simulations with first-principles machine learning potentials. Our simulations accurately reproduce the experimental phase diagram of carbon in the region around the GDL triple point and show that liquid carbon crystallises spontaneously upon cooling at constant pressure. Surprisingly, metastable graphite crystallises in the domain of diamond thermodynamic stability at pressures above the triple point. Furthermore, whereas diamond crystallises through a classical nucleation pathway, graphite follows a two-step process in which low-density fluctuations forego ordering. Calculations of the nucleation rates of the two competing phases confirm this result and reveal a manifestation of Ostwald’s step ruleOstwald (1897) where the strong metastability of graphite hinders the transformation to the stable diamond phase. Our results provide a new key to interpreting melting and recrystallisation experiments and shed light on nucleation kinetics in polymorphic materials with deep metastable states.
@@
The extremely diverse technological applications of carbon materials stem from the different types of covalent bonding with tetrahedral patterns in diamond and tetrahedral amorphous carbon on one hand, and hexagonal motifs in graphite or graphitic carbon on the other. Such differences in covalent bonding engender large metastability of both common crystalline phases, graphite, and diamond. These chemical features dictate the phase diagram of carbon, which exhibits a graphite-liquid-diamond triple point at about 12 GPa and 4500 K. Whereas the "diamond-in-the-sky" hypothesisRoss (1981) and recent experiments on nuclear fusion by laser shock compressionZylstra et al. (2022) have prompted several studies of the phase diagram of carbon and hydrocarbons at extreme pressures, the GDL triple point region has not been thoroughly addressed either by experiments or simulations. Conversely, understanding the thermodynamics and kinetics of phase changes in this region is of paramount importance for materials and planetary sciences. For example, in geosciences, diamonds reveal tectonic processes in Earth-like planets and the deep-earth cycle of carbon and other light elements.Timmerman et al. (2023) In materials processing, in turn, the liquid phase is an intermediate in the laser synthesis of carbon materials including artificial diamonds, nanodiamonds, and N-vacancy doped diamonds for quantum computing, and its crystallisation behaviour upon cooling determines the properties of the synthesis products.Hull, Raj, and Saykally (2020)
@@
Ostwald hypothesized that nucleation from the melt does not necessarily proceed directly to the thermodynamically most stable phase but through the phase that is separated by the lowest free energy barrier from the liquid.Ostwald (1897) This effect significantly accelerates crystallisation rates in the presence of metastable critical points.Talanquer and Oxtoby (1998) Molecular simulations of nucleation from the melt confirmed Ostwald’s step rule hypothesis in Lennard-Jones liquids, colloidal particles, and protein crystallisation.ten Wolde, Ruiz-Montero, and Frenkel (1995); Schilling et al. (2010); ten Wolde and Frenkel (1999) In the latter case, a two-step process was observed experimentally using dynamic light scattering and time-resolved spectroscopy. Non-classical multi-stage nucleation pathways have also been discussed in the context of mineralisation from solution, with abundant numerical and experimental evidence,Gebauer et al. (2014) and in the case of polymorphic materials the precipitation of metastable phases has been observed.Navrotsky (2004) Nevertheless, it is unexpected to encounter such a rich phenomenology in the homogeneous nucleation of a simple monoatomic system like carbon.
@@
In this Letter, we investigate the kinetic aspects of the crystallisation of diamond and graphite probing homogeneous nucleation from molten carbon at moderate pressures using machine-learning accelerated quantum-accurate molecular dynamics simulations. The range of pressures considered, between 5 and 30 GPa, corresponds to the pressure of the Earth’s mantle between 120 and 750 Km. Our simulations show the unexpected spontaneous crystallisation of metastable graphite at pressures up to 15 GPa, well within the domain of thermodynamic stability of diamond, and unravel the fundamentally distinct molecular mechanisms that lead to the nucleation of diamond and graphite. Our simulations exhibit a combination of these effects that are the consequence of Ostwald’s step rule and the large free energy barrier between graphite and diamond. These observations open new perspectives in the theory of crystallisation of single-component systems with metastable phases.
Results and discussion
@@
We first discuss the phase diagram of carbon in the region of the GDL triple point (Figure 1a). The triple point was initially determined at T$=4100\pm 200$ K, and P$=12.5\pm 1$ GPa by extrapolating the experimental graphite-diamond coexistence line to high temperatures.Kennedy and Kennedy (1976) Extrapolating the melting curves of graphiteTogaya (1997) and diamondBundy et al. (1996) suggests that the triple point should be at a higher temperature, between 4400 and 4800 K. However, in general, the experimental determination of the GDL triple point entails large uncertainties.Day (2012) In contrast to earlier measurements,Bundy (1963a) experiments now agree that the melting line of diamond has a positive slope while that of graphite has a negative slope at the triple point. The melting temperature of graphite also exhibits a maximum at $\sim$5 GPa. This was initially thought to be indicative of a liquid-liquid phase transition, which has not been confirmed by further simulations or experiments.Wu et al. (2002); Wang, Scandolo, and Car (2005)
@@
The graphite-liquid and diamond-liquid coexistence lines along with their metastable extensions were computed by large-scale phase-coexistence MD simulations. Our simulations successfully reproduce the experimental phase diagram and validate the accuracy of the machine learning potential utilized in this study. The intersection between the two melting curves establishes a GDL triple point at P=10.5$\pm$0.5 GPa and T=4650$\pm$50 K, at a slightly lower pressure than in experiments but well within the temperature range. The diamond-liquid coexistence line agrees with former FPMD simulations.Correa, Bonev, and Galli (2006); Wang, Scandolo, and Car (2005) Remarkably, our model reproduces the reentrant melting line of graphite in excellent agreement with measurements.Togaya (1997) This, to our knowledge, was never replicated in former molecular simulations. The reason for the negative slope of the melting line of graphite near the triple point is that the density of liquid carbon is higher than that of graphite at pressures higher than 2 GPa (Figure 1b). The density crossover at 2 GPa corresponds to the change of slope of the graphite melting line from positive to negative. This characteristic is extremely consequential for the crystallisation behaviour of graphite and diamond and is not captured by either empirical potentials or machine learning models fitted on different functionals.Ghiringhelli et al. (2008); Marchant, Karasulu, and Partay (2022); Marchant et al. (2023) Notably, using different parameterisations of NEP does not have a significant effect on the melting lines and the location of the triple point. The phase diagrams reported in Figure S1 share the same features as those in Figure 1.
Figure 1: (a) Phase diagram of carbon in the moderate pressure range 2-35 GPa. The melting temperatures of diamond (blue diamonds) and graphite (red squares) computed by phase coexistence molecular dynamics are represented by symbols with error bars. The purple dashed line is our theoretical estimate of the graphite/diamond phase boundary. The experimental phase diagram (brown line) combines results from Bundy et al., 1996; Togaya, 1997(black triangles). The temperatures at which we observed spontaneous crystallisation of diamond (blue) and graphite (red) are indicated with the respective symbols without error bars. The dashed black line indicates the maximum pressure at which graphite nucleates faster than diamond. The dashed-dotted lines indicate the homogeneous nucleation temperature for graphite (red) and diamond (blue), defined as the temperature at which crystal nucleation is faster than $10^{14}$ m-3s-1. (b) Density of diamond, graphite at 5000K, and liquid carbon at both 4000 and 5000 K at different pressures.
@@
Due to the small size of molecular models and the use of periodic boundary conditions, MD simulations customarily overestimate the stability of liquid phases below the melting point and the formation of amorphous phases by quenching from the melt. However, MD runs in which liquid carbon is quenched at the rate of 60 K/ns exhibit spontaneous crystallisation into either diamond or graphite depending on the pressure. The temperatures at which spontaneous crystallisation was observed are reported in the phase diagram (Figure 1a) and define the lowest-temperature boundary of the region of metastability of supercooled liquid carbon. This leads us to conclude that liquid carbon, as opposed to the other group IV elemental liquids like silicon and germanium, is not a good glass former at high pressure. Therefore, quenching from the melt at high pressure is not a viable route to obtain tetrahedral amorphous carbon. The temperature of spontaneous crystallisation of graphite decreases with pressure with a more accentuated negative slope than the phase coexistence line. Conversely, the temperature of spontaneous crystallisation of diamond increases with pressure but at a similar rate as the melting curve.
@@
Strikingly, we observe spontaneous crystallisation of graphite at pressures as high as 15 GPa, which is about 5 GPa higher than the pressure of the GDL triple point and 7 GPa higher than the graphite-diamond phase boundary at the crystallisation temperature. The occurrence of a metastable structure in the crystallisation pathway is predicted by Ostwald’s step rule when the metastable phase has a lower free energy barrier than the thermodynamically stable one.Ostwald (1897); Stranski and Totomanow (1933) This condition corresponds to a closer structural similarity between the nucleating metastable structure and the parent liquid. We observe that the highest pressure at which graphite crystallizes spontaneously corresponds to the condition where the density of the liquid is midway between that of graphite and that of diamond. (see Figure 1b). This observation suggests that the density in the liquid is an important driver of crystallisation. In turn, the change of preference for spontaneous crystallisation from graphite to diamond is not dictated by any abrupt changes in the structure of liquid carbon: both the pair correlation function and the angular distribution function undergo gradual changes as the pressure increases (Figure S2). In most cases, the metastable nucleus transforms spontaneously into the stable phase as it grows. In our simulations, however, graphite persists, as the activation free energy to transform graphite into diamond is high and crystallization happens too rapidly. The crystallisation of metastable crystalline phases from rapidly quenched liquids has been observed for minerals and alloys, especially when simple low-density structures compete with complex structures, but it was not formerly hypothesized for carbon. The pressure conditions between the graphite-diamond phase coexistence line and 15 GPa are not sufficient to overcome the high barrier to transforming graphite into diamond over the typical MD timescales.Khaliullin et al. (2011); Zhu et al. (2020) However, experimentally graphite transforms directly into diamond upon compression at 12.5 GPa above 3000 K,Bundy (1963b) thus suggesting that graphite would be a long-lived metastable intermediate state in the crystallisation of diamond.
@@
[…]
To validate the observed metastable crystallisation of graphite from liquid carbon, we computed the nucleation rates of graphite and diamond at moderate supercooling. These calculations require the use of an enhanced sampling technique with the possibility of biasing the nucleation of either phase. For this purpose, we used forward flux sampling molecular dynamics.Li et al. (2009); Li, Donadio, and Galli (2009) To identify the graphite and diamond nucleating crystallites we adopted local order parameters based on the spherical harmonic expansion of the atomic environment ‘$q_{l}$-dot".ten Wolde, Ruiz-Montero, and Frenkel (1996); Ghiringhelli et al. (2007) For diamond we use $l=6$ and for graphite $l=3$ with the additional constraint that atoms must be three-fold coordinated. Figure 2(a-c) shows the nucleation rates of graphite and diamond as a function of temperature up to 4100 K at $P=$13, 15, and 17 GPa. These calculations confirm a preference for graphite crystallisation at any temperature at 13 GPa. Nucleation rates for the two phases are similar at 15 GPa with a slight preference for graphite, whereas crystallisation of diamond is predominant at 17 GPa.
The rates and the estimates of nuclei critical size obtained by FFS allow us to fit a model based on classical nucleation theory (CNT). In CNT, the free energy of formation of a crystalline nucleus of size $N$ in a liquid is the result of a negative term that accounts for the difference in chemical potential between the solid and the liquid ($\Delta\mu$) and a positive term that accounts for the cost of creating solid/liquid interface. The CNT rate can be expressed as:
@@
In this expression, $\rho$ is the number density of the solid, and $\gamma_{LS}$ is the average interfacial tension between the liquid and the solid phase. $\Delta\mu$ is usually approximated as $\Delta\mu=\Delta H_{m}(T_{m}-T)/T$ where the enthalpy of melting ($\Delta H_{m}$) and the melting temperature ($T_{m}$) are calculated from MD simulations at solid/liquid coexistence conditions, leaving $A$ and $\gamma_{LS}$ as the only unknowns. Figure 2(a-c) shows that CNT (continuous lines) fits the FFS nucleation rates of both graphite and diamond. We note that the fitting of the FFS rate against CNT does not rely on the assumption of a spherical critical nucleus, but only requires the shape to remain unchanged at different temperatures. $\Delta G^{\star}$ obtained from this fit provides an estimate of the effective liquid/solid interfacial tension for graphite and diamond nuclei assumed spherical, where the former is $\gamma_{LS,G}=1.15$ J/m2 and the latter is $\gamma_{LS,D}\sim 2.0$ J/m2 at $P=15$ GPa. At an atomistic level, the high value of $\gamma_{LS,D}$ at moderate pressure can be explained by the stark difference in the bonding environment between the liquid, mostly three-fold coordinated, and the four-fold coordinated diamond crystallite.Ghiringhelli et al. (2007) The much smaller value of $\gamma_{LS,G}$ suggests that the direct nucleation of diamond can occur only when the chemical potential term is dominant, i.e. when the liquid is substantially more supercooled with respect to diamond than to graphite.
@@
To study unbiased nucleation pathways and obtain an estimate of the nucleation rates of graphite and diamond at deep supercooling,Yuhara et al. (2015) we have performed direct MD simulations at pressures of 15 GPa and 15.5 GPa and temperature of 3650 K. Averaging over 10 runs at each pressure, we computed a mean first passage time (MFPT) $\tau_{MFPT}=21.1\pm 2.8$ ns corresponding to a nucleation rate of 1.7±$\pm 0.2\times 10^{33}$ s-1m-3 for graphite at 15 GPa, and $\tau_{MFPT}=16.3\pm 5.4$ ns corresponding to 2.3±$\pm 0.7\times 10^{33}$ s-1m-3 for diamond at 15.5 GPa. The (MFPT) nucleation rate of graphite is found to match well the CNT model fitted through the FFS rates (Figure 2e), further confirming the validity of CNT and the applicability of FFS for graphite nucleation.
Figure 3: Nucleation pathways of graphite (top row) and diamond (bottom row) from direct molecular dynamics simulations at pressures of 15 and 15.5 GPa and temperature of 3650 K. Bonds are coloured according to the value of the local order parameter.

### v2

@@
Abstract
@@
Experimental challenges in determining the phase diagram of carbon at temperatures and pressures near the graphite-diamond-liquid triple point are often related to the persistence of metastable crystalline or glassy phases, superheated crystals, or supercooled liquids. A deeper understanding of the crystallisation kinetics of diamond and graphite is crucial for effectively interpreting the outcomes of these experiments. Here, we reveal the microscopic mechanisms of diamond and graphite nucleation from liquid carbon through molecular simulations with first-principles machine learning potentials. Our simulations accurately reproduce the experimental phase diagram of carbon near the triple point and show that liquid carbon crystallises spontaneously upon cooling. Metastable graphite crystallises in the domain of diamond thermodynamic stability at pressures above the triple point. Furthermore, whereas diamond crystallises through a classical nucleation pathway, graphite follows a two-step process in which low-density fluctuations forego ordering. Calculations of the nucleation rates of the two competing phases confirm this result and reveal a manifestation of Ostwald’s step rule, where the strong metastability of graphite hinders the transformation to the stable diamond phase. Our results provide a new key to interpreting melting and recrystallisation experiments and shed light on nucleation kinetics in polymorphic materials with deep metastable states.
Introduction
@@
The crystallisation of carbon from the melt under extreme conditions is highly relevant to Earth and planetary science,Timmerman et al. (2023); Ross (1981), materials manufacturing, and nuclear fusion research.Zylstra et al. (2022) The thermodynamic conditions near the graphite-diamond-liquid (GDL) triple point are especially of interest for geological and technological applications, but high-pressure flash heating experiments aiming to resolve this region of the phase diagram of carbon exhibit large discrepancies.Bundy et al. (1996); Togaya (1997); Zhao et al. (2013); Kondratyev and Rakhel (2019); Liang et al. (2019); Yang et al. (2023) The diverse technological applications of carbon materials stem from the different types of covalent bonding with tetrahedral patterns in diamond and tetrahedral amorphous carbon on one hand, and hexagonal motifs in graphite or graphitic carbon on the other. Such differences in covalent bonding engender large metastability of both common crystalline phases, graphite and diamond. These chemical features dictate the phase diagram of carbon, which exhibits a GDL triple point at about 12 GPa and 4500 K. Whereas the diamond-in-the-sky hypothesisRoss (1981) and recent experiments on nuclear fusion by laser shock compressionZylstra et al. (2022) have prompted several studies of the phase diagram of carbon and hydrocarbons at extreme pressures, the GDL triple point region has not been thoroughly addressed either by experiments or simulations. Conversely, understanding the thermodynamics and kinetics of phase changes in this region is of paramount importance for materials and planetary sciences. For example, in geosciences, diamonds reveal tectonic processes in Earth-like planets and the deep-earth cycle of carbon and other light elements.Timmerman et al. (2023) In materials processing, the liquid phase is an intermediate in the laser synthesis of carbon materials, including artificial diamonds, nanodiamonds, and N-vacancy doped diamonds for quantum computing, and its crystallisation behaviour upon cooling determines the properties of the synthesis products.Hull, Raj, and Saykally (2020)
@@
Ostwald hypothesized that nucleation from the melt does not necessarily proceed directly to the thermodynamically most stable phase but through the phase that is separated by the lowest free energy barrier from the liquid.Ostwald (1897) This effect significantly accelerates crystallisation rates in the presence of metastable critical points.Talanquer and Oxtoby (1998) Molecular simulations of nucleation from the melt confirmed Ostwald’s step rule hypothesis in Lennard-Jones liquids, colloidal particles, and protein crystallisation.ten Wolde, Ruiz-Montero, and Frenkel (1995); Schilling et al. (2010); ten Wolde and Frenkel (1999) In the latter case, a two-step process was observed experimentally using dynamic light scattering and time-resolved spectroscopy.Vorontsova, Maes, and Vekilov (2015) Non-classical multi-stage nucleation pathways have also been discussed in the context of mineralisation from solution, with abundant numerical and experimental evidence,Gebauer et al. (2014), and in the case of polymorphic materials, the precipitation of metastable phases has been observed.Navrotsky (2004) Nevertheless, it is unexpected to encounter such a rich phenomenology in the homogeneous nucleation of a simple monoatomic system like carbon.
@@
In this work, we investigate the kinetic aspects of diamond and graphite crystallisation, probing homogeneous nucleation from molten carbon at moderate pressures using machine-learning-accelerated, quantum-accurate molecular dynamics simulations. The range of pressures considered, between 5 and 30 GPa, corresponds to the pressure of the Earth’s mantle between 120 and 750 Km. Our simulations show the unexpected spontaneous crystallisation of metastable graphite at pressures up to 15 GPa, well within the domain of thermodynamic stability of diamond, and unravel the fundamentally distinct molecular mechanisms that lead to the nucleation of diamond and graphite. The crystallisation pathways exhibit a combination of these effects that are the consequence of Ostwald’s step rule and the large free energy barrier between graphite and diamond. These observations open new perspectives in the theory of crystallisation of single-component systems with metastable phases.
Results
@@
We first discuss the phase diagram of carbon in the region of the GDL triple point (Figure 1a). Experimentally, the triple point was initially determined at T$=4100\pm 200$ K, and P$=12.5\pm 1$ GPa by extrapolating the graphite-diamond coexistence line to high temperatures.Kennedy and Kennedy (1976) Extrapolating the melting curves of graphiteTogaya (1997) and diamondBundy et al. (1996) suggests that the triple point should be at a higher temperature, between 4400 and 4800 K. However, in general, the experimental determination of the GDL triple point entails large uncertainties.Day (2012) In contrast to earlier measurements,Bundy (1963a) experiments now agree that the melting line of diamond has a positive slope while that of graphite has a negative slope at the triple point. The melting temperature of graphite also exhibits a maximum at $\sim$5 GPa. This was initially thought to be indicative of a liquid-liquid phase transition,Glosli and Ree (1999) which was not confirmed by further simulations or experiments.Wu et al. (2002); Wang, Scandolo, and Car (2005) Recent experiments suggest that the melting point of graphite at 2 GPa and that of diamond at 15 GPa may be over 6000 K, significantly higher than in previous works, adding further uncertainty to the carbon phase diagram and the location of the GDL triple point.Kondratyev and Rakhel (2019); Liang et al. (2019)
Here we compute the graphite-liquid and diamond-liquid coexistence lines, along with their metastable extensions, by large-scale phase-coexistence MD simulations using feed-forward neural network potentials fitted with an evolutionary strategy (neuroevolution potential, NEP)Fan et al. (2021) and trained on energies and forces computed at the level of density functional theory (DFT) either in the local density approximation (LDA)Deringer and Csányi (2017) and with a generalized gradient approximation functional with non-local van der Waals correlation ‘OptB88-vdW’.Klimeš, Bowler, and Michaelides (2010); Rowe et al. (2020) We dub these two models NEP@LDA and NEP@OptB88-vdW.
@@
Our simulations closely reproduce the most commonly accepted experimental phase diagram regardless of the chosen NEP model.Bundy et al. (1996); Savvatimskiy (2015) In the NEP@LDA diagram, the intersection between the two melting curves establishes a GDL triple point at P=10.5$\pm$0.5 GPa and T=4650$\pm$50 K, at a slightly lower pressure than in most experiments but well within the temperature range. Starting from the triple point, we computed the graphite-diamond coexistence line using the Gibbs-Duhem integration method.Kofke (1993) The diamond-liquid coexistence line agrees with former FPMD simulations.Correa, Bonev, and Galli (2006); Wang, Scandolo, and Car (2005) Remarkably, our model reproduces the reentrant melting line of graphite in excellent agreement with measurements.Togaya (1997) This, to our knowledge, was never replicated in former molecular simulations. The reason for the negative slope of the melting line of graphite near the triple point is that the density of liquid carbon is higher than that of graphite at pressures higher than 2 GPa (Figure 1b). The density crossover at 2 GPa corresponds to the change of slope of the graphite melting line from positive to negative. This characteristic is extremely consequential for the crystallisation behaviour of graphite and diamond and is rarely captured by either empirical potentials or machine learning models,Ghiringhelli et al. (2008) with few notable exceptions, such as the revised long-range carbon bond-order potential (LCBOPII) and the environment-dependent interatomic potential (EDIP).Colonna et al. (2009); Marks (2000); Marchant et al. (2023) Notably, using different parameterisations of NEP does not have a significant effect on the melting lines and the location of the triple point. The phase diagrams reported in Figure S1 share the same features as those in Figure 1.
Since the approximations in the LDA exchange and correlation functional entail inaccuracies in the equations of state and the phase diagram of materials, usually overestimating the density of both solids and liquids,French and Mattsson (2014) we repeated the calculation of the phase diagram of carbon with NEP@OptB88-vdW.Klimeš, Bowler, and Michaelides (2010); Rowe et al. (2020); Fan et al. (2024) The NEP@LDA and NEP@OptB88-vdW phase diagrams are reported side by side in Figure S2. The GDL triple point with the NEP@OptB88-vdW model is at the same temperature as NEP@LDA but at a higher pressure (P=17 GPa and T=4660 K). In general, the two phase diagrams exhibit the same features with just a pressure offset of $\sim 6.5$ GPa.
As expected from previous DFT calculations,French and Mattsson (2014) the equations of state of liquid carbon, computed at 5000 and 7000 K differ with NEP@LDA overestimating and NEP@OptB88-vdW underestimating the density compared to experiments (Figure S3).Kondratyev and Rakhel (2019) However, the structure of the liquid, characterised through radial distribution functions (Figure S4) and the analysis of the local coordination (Figure S5), exhibits the same behaviour as the pressure is increased. At any pressure between 5 and 30 GPa, compared to the NEP-OptB88-vdW liquid, the NEP@LDA liquid exhibits a larger number of tetrahedrally coordinated carbons that increases linearly with the pressure, and a lower number of twofold coordinated carbons that decreases. The number of threefold coordinated carbons has a maximum between 5 and 10 GPa for NEP@LDA and above 15 GPa for NEP@OptB88-vdW. The differences in the properties of the liquids obtained with the two models may be reconciled with a pressure shift of about 10 GPa. The liquid structure is consistent with previous studies based on DFT or accurate empirical potentials.Ghiringhelli et al. (2005, 2008); Colonna et al. (2009)
@@
Due to the small size of molecular models and the use of periodic boundary conditions, MD simulations customarily overestimate the stability of liquid phases below the melting point and the formation of amorphous phases by quenching from the melt. However, MD runs in which liquid carbon models equilibrated at 5000 K are quenched at the rate of 60 K/ns exhibit spontaneous crystallisation into either diamond or graphite, depending on the pressure. The temperatures at which spontaneous crystallisation was observed are reported in the phase diagram (Figure 1a and S2) and define the lowest-temperature boundary of the region of metastability of supercooled liquid carbon. This leads us to conclude that liquid carbon, as opposed to other group IV elemental liquids like silicon and germanium, is not a good glass former at high pressure. Therefore, quenching from the melt at high pressure is not a viable route to obtain tetrahedral amorphous carbon. The temperature of spontaneous crystallisation of graphite decreases with pressure with a more accentuated negative slope than the phase coexistence line. Conversely, the temperature of spontaneous crystallisation of diamond increases with pressure but with a similar slope as the melting curve.
@@
Strikingly, we observe spontaneous crystallisation of graphite at pressures as high as 15 GPa, which is about 5 GPa higher than the pressure of the GDL triple point and 7 GPa higher than the graphite-diamond phase boundary at the crystallisation temperature. The occurrence of a metastable structure in the crystallisation pathway is predicted by Ostwald’s step rule when the metastable phase has a lower free energy barrier than the thermodynamically stable one.Ostwald (1897); Stranski and Totomanow (1933) This condition corresponds to a closer structural similarity between the nucleating metastable structure and the parent liquid. We observe that the highest pressure at which graphite crystallizes spontaneously corresponds to the condition where the density of the liquid is midway between that of graphite and that of diamond (see Figure 1b). This observation suggests that the density of the liquid is an important driver of crystallisation. In turn, the change of preference for spontaneous crystallisation from graphite to diamond is not dictated by any abrupt changes in the structure of liquid carbon: the equation of state, pair correlation function, and local coordination undergo gradual changes as the pressure increases (Figures S3-S5) in accordance with previous studies.Ghiringhelli et al. (2005); Colonna et al. (2009) It is worth noting that at these pressures, the liquid is still predominantly three-fold coordinated. In most cases, the metastable nucleus transforms spontaneously into the stable phase as it grows. In our simulations, however, graphite persists, as the activation free energy to transform graphite into diamond is high, and crystallisation happens too rapidly. The crystallisation of metastable crystalline phases from rapidly quenched liquids has been observed for minerals and alloys, especially when simple low-density structures compete with complex structures, but it was not formerly hypothesised for carbon. The pressure conditions between the graphite-diamond phase coexistence line and 15 GPa are insufficient to overcome the high barrier to transforming graphite into diamond over the typical MD timescales.Khaliullin et al. (2011); Zhu et al. (2020) However, experimentally graphite transforms directly into diamond upon compression at 12.5 GPa above 3000 K,Bundy (1963b) thus suggesting that graphite would be a long-lived metastable intermediate state in the crystallisation of diamond.
Spontaneous crystallisation of metastable graphite above the graphite-diamond coexistence line also occurs in simulations performed with the NEP@OptB88-vdW model (Figure S2). In this case, we observed metastable nucleation of graphite up to 19 GPa at 3670 K, where the corresponding graphite-diamond phase boundary is at 14.5 GPa (Figure S2). Since the two NEP models provide the same qualitative physical picture of the thermodynamics of carbon in the region of interest and of the kinetics of crystallisation, we restrict the analysis in the following sections to results obtained using the NEP@LDA model.
@@
To validate the observed metastable crystallisation of graphite from liquid carbon, we computed the nucleation rates of graphite and diamond at moderate supercooling. These calculations require the use of an enhanced sampling technique with the possibility of biasing the nucleation of either phase. For this purpose, we used forward flux sampling molecular dynamics.Li et al. (2009); Li, Donadio, and Galli (2009) To identify the graphite and diamond nucleating crystallites, we adopted local order parameters based on the spherical harmonic expansion of the atomic environment ‘$q_{l}$-dot".ten Wolde, Ruiz-Montero, and Frenkel (1996); Ghiringhelli et al. (2007) For diamond, we use $l=6$ and for graphite $l=3$ with the additional constraint that atoms must be three-fold coordinated. Figure 2(d-f) shows the nucleation rates of graphite and diamond as a function of temperature up to 4100 K at $P=$13, 15, and 17 GPa. These calculations confirm a preference for graphite crystallisation at any temperature at 13 GPa. Nucleation rates for the two phases are similar at 15 GPa with a slight preference for graphite, whereas crystallisation of diamond is predominant at 17 GPa.
The rates and the estimates of the critical size of the nuclei obtained by FFS allow us to fit a model based on classical nucleation theory (CNT). In CNT, the free energy of formation of a crystalline nucleus of size $N$ in a liquid is the result of a negative term that accounts for the difference in chemical potential between the solid and the liquid ($\Delta\mu$) and a positive term that accounts for the cost of creating solid/liquid interface. The CNT rate can be expressed as:
@@
In this expression, $\rho$ is the number density of the solid, and $\gamma_{LS}$ is the average interfacial tension between the liquid and the solid phase. $\Delta\mu$ is usually approximated as $\Delta\mu=\Delta H_{m}(T_{m}-T)/T$ where the enthalpy of melting ($\Delta H_{m}$) and the melting temperature ($T_{m}$) are calculated from MD simulations at solid/liquid coexistence conditions, leaving $A$ and $\gamma_{LS}$ as the only unknowns. Figure 2(d-f) shows that CNT (continuous lines) fits the FFS nucleation rates of both graphite and diamond. We note that the fitting of the FFS rate against CNT does not rely on the assumption of a spherical critical nucleus, but only requires the shape to remain unchanged at different temperatures. $\Delta G^{\star}$ obtained from this fit provides an estimate of the effective liquid/solid interfacial tension for graphite and diamond nuclei assumed spherical, where the former is $\gamma_{LS,G}=1.15$ J/m2 and the latter is $\gamma_{LS,D}\sim 2.0$ J/m2 at $P=15$ GPa. At an atomistic level, the high value of $\gamma_{LS,D}$ at moderate pressure can be explained by the stark difference in the bonding environment between the liquid, mostly three-fold coordinated, and the four-fold coordinated diamond crystallite.Ghiringhelli et al. (2007) The much smaller value of $\gamma_{LS,G}$ suggests that the direct nucleation of diamond can occur only when the chemical potential term is dominant, i.e., when the liquid is substantially more supercooled with respect to diamond than to graphite.
@@
To study unbiased nucleation pathways and obtain an estimate of the nucleation rates of graphite and diamond at deep supercooling,Yuhara et al. (2015) we have performed direct MD simulations at pressures of 15 GPa and 15.5 GPa and a temperature of 3650 K. Averaging over 10 runs at each pressure, we computed a mean first passage time (MFPT) $\tau_{MFPT}=21.1\pm 2.8$ ns corresponding to a nucleation rate of 1.7±$\pm 0.2\times 10^{33}$ s-1m-3 for graphite at 15 GPa, and $\tau_{MFPT}=16.3\pm 5.4$ ns corresponding to 2.3±$\pm 0.7\times 10^{33}$ s-1m-3 for diamond at 15.5 GPa. The (MFPT) nucleation rate of graphite is found to match well the CNT model fitted through the FFS rates (Figure 2e), further confirming the validity of CNT and the applicability of FFS for graphite nucleation.